### 20 Mar 2024 v2

# JEL Codes: C21, C31, C57

Abstract. In settings with interference, researchers commonly define estimands using exposure mappings to summarize neighborhood variation in treatment assignments. This paper studies the causal interpretation of these estimands under weak restrictions on interference. We demonstrate that the estimands can exhibit unpalatable sign reversals under conventional identification conditions. This motivates the formulation of sign preservation criteria for causal interpretability. To satisfy preferred criteria, it is necessary to impose restrictions on interference, either in potential outcomes or selection into treatment. We provide sufficient conditions and show that they can be satisfied by nonparametric models with interference in both the outcome and selection stages.
JEL Codes: C21, C31, C57
Keywords: causal inference, interference, peer effects
1 Introduction
Consider a set of $n$ units, and for each unit $i$, let $Y_{i}\in\mathbb{R}$ denote its outcome and $D_{i}$ an indicator for whether the unit is treated. We study settings with interference where outcomes may depend on the entire treatment assignment vector $\bm{D}=(D_{i})_{i=1}^{n}$. Since the causal effect of $\bm{D}$ on $Y_{i}$ is difficult to convey on account of the dimensionality of $\bm{D}$, the literature commonly employs estimands defined by exposure mappings to parsimoniously summarize key dimensions of the effect (Aronow and Samii, 2017; Manski, 2013).
For some function $f\colon\{1,\ldots n\}\times\{0,1\}^{n}\rightarrow\mathbb{R}^{d_{t}}$ chosen by the econometrician, define for each $i$ its exposure mapping $T_{i}=f(i,\bm{D})$. We study the estimand
$\tau(t,t^{\prime})=\frac{1}{n}\sum_{i=1}^{n}{\bf E}\big{[}{\bf E}[Y_{i}\mid T_{i}=t,\mathcal{C}_{i}]-{\bf E}[Y_{i}\mid T_{i}=t^{\prime},\mathcal{C}_{i}]\big{]}\quad\text{for }t,t^{\prime}\in\mathbb{R}^{d_{t}},$ (1)
where $\mathcal{C}_{i}$ is a set of control variables for unit $i$. Variants of $\tau(t,t^{\prime})$ are widely used in the literature (e.g. Forastiere et al., 2021; Leung, 2022; Ogburn et al., 2022). For example, suppose $T_{i}=(D_{i},N_{i})$ where $N_{i}$ is the number of treated units within some given distance of $i$. Then for $t=(1,0)$ and $t^{\prime}=(0,0)$, $\tau(t,t^{\prime})$ measures the direct “effect” of the treatment for units with no treated neighbors, while for $t=(0,2)$ and $t^{\prime}=(0,1)$, it measures an indirect or spillover “effect.”
We put “effect” in quotations because the causal interpretation of $\tau(t,t^{\prime})$ is a matter of some controversy (Auerbach et al., 2024; Sävje, 2024b). Studying an experimental setting, Sävje (2024a) refers to $\tau(t,t^{\prime})$ as the “expected exposure effect,” suggesting the view that it has a causal interpretation. Harshaw et al. (2022) propose a larger class of estimands that includes $\tau(t,t^{\prime})$ as a special case, and they explicitly label these causal estimands. The latter two references informally discuss the interpretation of $\tau(t,t^{\prime})$, but missing from the literature is a rigorous study of its causal content. This paper aims to fill this gap. [^1] The first working paper version of Leung (2022) provides some formal discussion for the case of Bernoulli randomized designs (Leung, 2019, §A.1).
Heckman and Pinto (2023) distinguish between statistical and ceteris peribus comparisons, noting that only the latter constitute causal effects. The estimand $\tau(t,t^{\prime})$ is a statistical comparison, so to formally establish that it has causal content, one must first define the ceteris peribus comparisons of interest and then prove an identification result demonstrating that $\tau(t,t^{\prime})$ is, in some sense, informative for the comparisons. This requires a rigorous definition of what it means for an estimand to be “informative for” ceteris peribus comparisons, and we adopt sign preservation criteria for this purpose (Bugni et al., 2023).
We study a collection of widely used exposure mappings in the literature, each of which defines an aggregate statistical comparison $\tau(t,t^{\prime})$. We define the ceteris peribus comparisons of interest to be partial effects that constitute unit-level analogs of the aggregate comparison. We then present simple examples demonstrating that the sign of $\tau(t,t^{\prime})$ can be entirely inconsistent with those of the unit-level effects even under standard identification conditions. For instance, in § 2, we consider the population difference in means, which corresponds to $\tau(1,0)$ with $T_{i}=D_{i}$, and take the ceteris peribus comparisons of interest to be the unit-level treatment effects $Y_{i}(1,\bm{d}_{-i})-Y_{i}(0,\bm{d}_{-i})$, where $\bm{d}_{-i}$ is the subvector of $\bm{d}\in\{0,1\}^{n}$ excluding $i$’s assignment. We show that unit-level treatment effects can be homogeneous, yet $\tau(1,0)$ can have the opposite sign under a completely randomized design.
These examples motivate the definition of sign preservation criteria, which state that, if all unit-level partial effects have identical signs, then $\tau(t,t^{\prime})$ has the same sign. We discuss several generic criteria and make the point that the relevant criterion for a given problem depends on the specific choice of exposure mapping and ceteris peribus comparisons. A criterion may rule out unpalatable sign reversals for one choice yet fail to do so for another.
We prove new identification results that provide interpretable sufficient conditions under which $\tau(t,t^{\prime})$ satisfies sign preservation criteria. These conditions primarily involve restrictions on interference. In observational settings, we may anticipate interference not only in potential outcomes but also in selection into treatment (Balat and Han, 2023). For instance, suppose $D_{i}$ is an indicator for whether the individual smokes and $Y_{i}$ is a health outcome. Second-hand smoke induces interference in potential outcomes, and peer effects in smoking adoption induce interference in treatment selection. Our analysis demonstrates that sign preservation generally requires restrictions on interference either for the outcome or the treatment selection stage. This is similar to how the Wald estimand for instrumental variables obtains a causal interpretation under restrictions on heterogeneity either for potential outcomes (homogeneous treatment effects) or treatment selection (monotonicity; Imbens and Angrist, 1994; Vytlacil, 2002).
There is a great deal of recent work on the causal interpretations of common regression estimands in the absence of interference when treatment effects are heterogeneous (e.g. Blandhol et al., 2022; Bugni et al., 2023; Chetverikov et al., 2023; de Chaisemartin and d’Haultfoeuille, 2020; Goldsmith-Pinkham et al., 2024, and references therein). We study nonparametric estimands under interference, and our examples illustrate how the main impediment to causal interpretability is not heterogeneity but rather the combination of interference in potential outcomes and selection into treatment.
The next section introduces some of the main themes by focusing on the simple case of the population difference in means under interference. We subsequently generalize the analysis to the larger class of $K$-neighborhood exposure mappings. In § 3 and § 4, we present our main results, which provide conditions under which $\tau(t,t^{\prime})$ satisfies different sign preservation criteria. Finally, § 5 concludes.
2 Difference in Means
This section considers the case $T_{i}=D_{i}$ and $\mathcal{C}_{i}=\varnothing$, so that
$\tau(1,0)=\frac{1}{n}\sum_{i=1}^{n}{\bf E}\big{[}{\bf E}[Y_{i}\mid D_{i}=1]-{\bf E}[Y_{i}\mid D_{i}=0]\big{]},$
the population difference in means. For all $i$, let $Y_{i}(\cdot)$ be a potentially random mapping from $\{0,1\}^{n}$ to $\mathbb{R}$. We interpret $Y_{i}(\bm{d})$ as the potential outcome of unit $i$ under the counterfactual that the treatment assignment vector is $\bm{d}\in\{0,1\}^{n}$ so that $Y_{i}=Y_{i}(\bm{D})$. [^2] Throughout the paper, $\bm{D}$ is considered random, while potential outcomes and controls may either be random or fixed. Suppose treatment assignment is as good as random in that
$Y_{i}(\cdot)\perp\!\!\!\perp\bm{D}\quad\text{for all }i.$ (2)
The discussion that follows relates to criticisms of the use of difference in means in analyses of vaccine trials (Eck et al., 2018; VanderWeele and Tchetgen, 2011). We provide a complementary perspective in terms of sign preservation criteria, but the main purpose of this section is to provide a simple illustration of ideas that will be generalized to a larger class of exposure mappings and observational settings in § 3.
2.1 Sign Reversal and Preservation
Because $\tau(1,0)$ compares treated and untreated units in the aggregate, we define the analogous unit-level partial effects as the ceteris peribus comparisons of interest. For $d\in\{0,1\}$ and $\bm{d}\in\{0,1\}^{n}$, let $Y_{i}(d,\bm{d}_{-i})$ denote $i$’s potential outcome under the treatment assignment vector that swaps out the $i$th component of $\bm{d}$ for $d$. For any $i$, define the comparison set
$\mathcal{T}_{i}^{*}(1,0)=\big{\{}Y_{i}(1,\bm{d}_{-i})-Y_{i}(0,\bm{d}_{-i})\colon\bm{d}_{-i}\in\{0,1\}^{n-1}\big{\}}.$
We next present an example under which the sign of $\tau(1,0)$ is entirely inconsistent with the signs of elements in the comparison set for all units $i$ despite (2) holding.
Suppose $n=2$, so that we may write $Y_{i}(\bm{d})=Y_{i}(d_{1},d_{2})$ where $d_{1}$ is unit 1’s counterfactual assignment and $d_{2}$ is unit 2’s. Let
$Y_{1}(0,0)=Y_{2}(0,0)=0$ ${\bf P}(\bm{D}=(0,0)\mid D_{1}=0)={\bf P}(\bm{D}=(0,0)\mid D_{2}=0)=p_{1}$
$Y_{1}(1,0)=Y_{2}(0,1)=1$ ${\bf P}(\bm{D}=(1,0)\mid D_{1}=1)={\bf P}(\bm{D}=(0,1)\mid D_{2}=1)=p_{2}$
$Y_{1}(0,1)=Y_{2}(1,0)=2$ ${\bf P}(\bm{D}=(0,1)\mid D_{1}=0)={\bf P}(\bm{D}=(1,0)\mid D_{2}=0)=p_{3}$
$Y_{1}(1,1)=Y_{2}(1,1)=3$ ${\bf P}(\bm{D}=(1,1)\mid D_{1}=1)={\bf P}(\bm{D}=(1,1)\mid D_{2}=1)=p_{4}$
The restriction to $n=2$ is for simplicity, and the example is easily scaled up by considering a large sample of identical dyads. Notice there is no heterogeneity across the two units, so
$\tau(1,0)={\bf E}[Y_{1}\mid D_{1}=1]-{\bf E}[Y_{1}\mid D_{1}=0]=3p_{4}+p_{2}-2p_{3}.$
Observe that all unit-level treatment effects are positive: $Y_{1}(1,d_{2})-Y_{1}(0,d_{2})=Y_{2}(d_{1},1)-Y_{2}(d_{1},0)=1$ for any $d_{1},d_{2}\in\{0,1\}$. Yet it is straightforward to construct assignment mechanisms such that $\tau(1,0)<0$. For example, consider the simple case of complete randomization with half of units (one in each dyad) allocated to treatment: ${\bf P}(\bm{D}=(1,0))={\bf P}(\bm{D}=(0,1))=0.5$. Then $(p_{1},p_{2},p_{3},p_{4})=(0,1,1,0)$, so $\tau(1,0)=-1$, the opposite of the unit-level effects. Consequently, $\tau(1,0)$ is not informative for the ceteris peribus comparisons even under randomized assignment.
The sign reversal occurs for two reasons. First, since there is interference in potential outcomes, $D_{i}$ does not fully summarize the effect of $\bm{D}$ on $Y_{i}$, so $T_{i}=D_{i}$ is what Sävje (2024a) calls a “misspecified” exposure mapping. Second, treatment assignments are dependent across units. Here this is due to complete randomization, but in observational settings, dependence may arise due to interference in selection into treatment.
This example motivates the following sign preservation criterion, which formalizes a minimal sense in which $\tau(t,t^{\prime})$ is informative for elements in the comparison set.
Definition 1.
$\tau(1,0)$ satisfies treatment sign preservation if $\min_{i}\min\mathcal{T}_{i}^{*}(1,0)\geq 0$ implies $\tau(1,0)\geq 0$ and $\max_{i}\max\mathcal{T}_{i}^{*}(1,0)\leq 0$ implies $\tau(1,0)\leq 0$.
That is, $\tau(1,0)$ is positive (negative) if all unit-level treatment effects are positive (negative).
In the previous example, treatment sign preservation is violated under randomized assignment (2), which demonstrates that additional restrictions are required beyond conventional identification conditions. For example, under the stable unit treatment value assumption (SUTVA), the exposure mapping is no longer misspecified, and $\tau(1,0)$ equals the usual average treatment effect. Alternatively, if treatments are Bernoulli randomized, then for $\bm{D}=(D_{i},\bm{D}_{-i})$, we have $D_{i}\perp\!\!\!\perp\bm{D}_{-i}$, so
$\tau(1,0)=\frac{1}{n}\sum_{i=1}^{n}{\bf E}\big{[}Y_{i}(1,\bm{D}_{-i})-Y_{i}(0,\bm{D}_{-i})\big{]}.$
In either case, $\tau(1,0)$ equals a quantity that trivially satisfies Definition 1. These examples are well understood in the literature, but the point we wish to emphasize is that the first case is a restriction on interference for potential outcomes while the second is a restriction on dependence in the assignment mechanism. We will see how these generalize to observational settings in the next section.
2.2 General Sign Preservation
Through the course of the paper, we will define several generic sign preservation criteria and argue that which is preferable depends on the exposure mapping and ceteris peribus comparisons of interest. To provide a basic illustration of this point, we compare treatment sign preservation with a criterion proposed by Sävje (2024b), which has the advantage of applying to arbitrary exposure mappings. However, when specialized to the case of $T_{i}=D_{i}$, it becomes apparent that the criterion is too weak to rule out the sign reversal in § 2.1. Therefore, it is not the relevant criterion, provided that the ceteris peribus comparisons of interest are given by $\mathcal{T}_{i}^{*}(1,0)$. However, interest may center on other comparisons, which we discuss in § 2.3 below.
Definition 2 (Sävje (2024b) Definition 1).
For $t,t^{\prime}\in\mathbb{R}^{d_{t}}$ and an exposure mapping $f$, define the comparison set
$\mathcal{T}_{i}(t,t^{\prime})=\big{\{}Y_{i}(\bm{d})-Y_{i}(\bm{d}^{\prime})\colon\bm{d},\bm{d}^{\prime}\in\{0,1\}^{n},f(i,\bm{d})=t,f(i,\bm{d}^{\prime})=t^{\prime}\big{\}}.$
The estimand $\tau(t,t^{\prime})$ defined by exposure mapping $f$ satisfies general sign preservation if $\min_{i}\min\mathcal{T}_{i}(t,t^{\prime})\geq 0$ implies $\tau(t,t^{\prime})\geq 0$ and $\max_{i}\max\mathcal{T}_{i}(t,t^{\prime})\leq 0$ implies $\tau(t,t^{\prime})\leq 0$.
That is, if all assignment vectors that induce an exposure mapping of $t$ result in larger (smaller) potential outcomes than those of all assignment vectors that induce $t^{\prime}$, then $\tau(t,t^{\prime})$ is positive (negative). In the § 2.1 example, neither $\min_{i}\min\mathcal{T}_{i}(1,0)\geq 0$ nor $\max_{i}\max\mathcal{T}_{i}(1,0)\leq 0$ since $Y_{1}(1,1)-Y_{1}(0,1)>0$ and $Y_{1}(1,0)-Y_{1}(0,1)<0$, so general sign preservation imposes no restrictions on the sign of $\tau(1,0)$.
Notice that when $T_{i}=D_{i}$, $t=1$, and $t^{\prime}=0$,
$\mathcal{T}_{i}(1,0)=\left\{Y_{i}(1,\bm{d}_{-i})-Y_{i}(0,\bm{d}_{-i}^{\prime})\colon\bm{d},\bm{d}^{\prime}\in\{0,1\}^{n}\right\}.$
Whereas our comparison set $\mathcal{T}_{i}^{*}(1,0)$ contains only unit-level treatment effects since $\bm{d}_{-i}$ is held constant, $\mathcal{T}_{i}(1,0)$ contains many additional comparisons. In the § 2.1 example, the latter includes the comparison $Y_{1}(1,0)-Y_{1}(0,1)$, but the former does not because it is not a ceteris peribus effect of treatment.
The difference in comparison sets has very different implications for the attainability of sign preservation. Proposition 1 of Sävje (2024b) shows that $\tau(t,t^{\prime})$ satisfies general sign preservation with no further assumptions (other than (2)). The example in § 2.1 shows that treatment sign preservation requires additional restrictions because it is a stronger criterion.
2.3 Total and Overall Effects
Our reading of $\tau(1,0)$ is that it is intended to summarize unit-level partial effects, which involve isolating variation in the ego’s treatment while holding fixed treatments assigned to alters. This motivates our definition of the comparison set $\mathcal{T}_{i}^{*}(1,0)$, which only contains unit-level partial effects. However, one might instead consider total effects (Bugni et al., 2023, Remark 3.1). Sävje (2024a) refers to $\tau(1,0)$ as the “average distributional shift effect” (ADSE), writing, “Because the two conditional distributions [of $\bm{D}$], corresponding to $D_{i}=1$ and $D_{i}=0$, potentially are different, the ADSE estimand can capture both the direct effect of $D_{i}$ on the outcome of unit $i$ and effects due to the distributional shift of the marginalization” (p. 25 of his supplementary appendix).
Giving “distributional shift effect” causal meaning requires specifying a ceteris peribus comparison of interest. One attempt might be $Y_{1}(1,D_{2}(1))-Y_{1}(0,D_{2}(0))$ (in an $n=2$ example) where $D_{2}(d)$ is the counterfactual treatment of unit 2 when unit 1’s assignment is fixed at $d$. This is a total effect that combines the direct effect of manipulating unit 1’s assignment with a mediated effect through unit 2’s assignment (Pearl, 2022). A mechanism for the latter might be peer effects in selection into treatment, and $D_{2}(1)-D_{2}(0)$ might be thought of as a partial equilibrium effect. Because we doubt that $\tau(t,t^{\prime})$ can be informative for unit-level total effects of this type in our setup without extremely restrictive assumptions, we limit our attention to partial effects.
In the biostatistics literature there is the notion of an “overall effect,” which is $\tau(t,t^{\prime})$ when $T_{i}$ is the overall fraction of treated units (Hudgens and Halloran, 2008; VanderWeele and Tchetgen, 2011). An example given by Sävje (2024b) is a comparison of two vaccine policies respectively treating 75 and 25 percent of the population. Still, $\tau(t,t^{\prime})$ remains a statistical comparison, and one can construct sign reversal examples similar to those in this paper in which $Y_{i}(\bm{d})$ is weakly increasing in each component of $\bm{d}$ for all $i$ (i.e. everyone benefits from more vaccinated peers), yet $\tau(0.75,0.25)<0$.
This is to emphasize that, whether it is the difference in means, overall effect, or any other variation of $\tau(t,t^{\prime})$, these statistical comparisons are generally informative for ceteris peribus comparisons only under further restrictions on potential outcomes or the assignment mechanism beyond a standard identification condition such as (2). For instance, much of the literature on the overall effect additionally assumes partial interference and Bernoulli or complete randomization within clusters, in which case the causal interpretation of the overall effect is well understood (Hudgens and Halloran, 2008).
3 K-Neighborhood Exposure Mappings
We next consider observational settings and a broader class of exposure mappings. Instead of randomized assignment (2), we maintain the following unconfoundedness condition.
Assumption 1 (Unconfoundedness).
$Y_{i}(\cdot)\perp\!\!\!\perp\bm{D}\mid\mathcal{C}_{i}$ for all $i$.
Leung and Loupos (2024) study the economic content of this assumption and choice of controls $\mathcal{C}_{i}$. We provide further discussion below.
Our results pertain to the following mild generalization of (1):
$\tau(t,t^{\prime})=\frac{1}{m_{n}}\sum_{i\in\mathcal{M}_{n}}\big{(}{\bf E}[Y_{i}\mid T_{i}=t,\mathcal{C}_{i}]-{\bf E}[Y_{i}\mid T_{i}=t^{\prime},\mathcal{C}_{i}]\big{)}$
for $\mathcal{M}_{n}\subseteq\{1,\ldots,n\}$ and $m_{n}=\lvert\mathcal{M}_{n}\rvert$. Depending on the exposure mapping, restricting the comparison to a subpopulation $\mathcal{M}_{n}$ may be necessary for overlap, ensuring that both events $T_{i}=t$ and $T_{i}=t^{\prime}$ occur with nontrivial probability. We provide examples below.
3.1 Exposure Mappings
Suppose units are connected through a nonrandom, unweighted network $\bm{A}$ with no self-links. Let $\mathcal{N}(i,K)$ denote unit $i$’s $K$-neighborhood, the subset of units at most path distance $K$ from $i$ in $\bm{A}$, noting that $\mathcal{N}(i,0)=\{i\}$. For any $\bm{d}\in\{0,1\}^{n}$, let $\bm{d}_{\mathcal{N}(i,K)}$ be the subvector restricted to units in $\mathcal{N}(i,K)$ and $\bm{d}_{-\mathcal{N}(i,K)}$ the subvector restricted to $\{1,\ldots,n\}\backslash\mathcal{N}(i,K)$. It will often be convenient to partition $\bm{d}$ as $(\bm{d}_{\mathcal{N}(i,K)},\bm{d}_{-\mathcal{N}(i,K)})$ and write $Y_{i}(\bm{d}_{\mathcal{N}(i,K)},\bm{d}_{-\mathcal{N}(i,K)})\equiv Y_{i}(\bm{d})$.
In the remainder of the paper, we focus on the class of $K$-neighborhood exposure mappings where $f(i,\bm{d})=f(i,\bm{d}^{\prime})$ for all $\bm{d},\bm{d}^{\prime}\in\{0,1\}^{n}$ such that $\bm{d}_{\mathcal{N}(i,K)}=\bm{d}_{\mathcal{N}(i,K)}^{\prime}$. In other words, the exposure mapping only depends on treatments assigned to the ego’s $K$-neighborhood, where $K$ is usually chosen much smaller than the typical path distance between units. Abusing notation, we write
$f(\bm{d}_{\mathcal{N}(i,K)})\equiv f(i,\bm{d}),$
so that $f(\bm{D}_{\mathcal{N}(i,K)})=T_{i}$.
Most of the literature assumes that the exposure mapping is correctly specified in that $Y_{i}(\bm{d})=Y_{i}(\bm{d}^{\prime})$ for all $\bm{d},\bm{d}^{\prime}\in\{0,1\}^{n}$ such that $f(\bm{d}_{\mathcal{N}(i,K)})=f(\bm{d}_{\mathcal{N}(i,K)}^{\prime})$ (Sävje, 2024a). In other words, $f$ fully summarizes interference, so $Y_{i}$ only depends on $\bm{d}$ through $f(\bm{d}_{\mathcal{N}(i,K)})$. In this case, $\tau(t,t^{\prime})$ is well known to have a transparent causal interpretation under an analog of Assumption 1 (Aronow and Samii, 2017; Forastiere et al., 2021; Ogburn et al., 2022).
Sävje (2024a) argues that $T_{i}$ is usually misspecified in practice, and Leung (2022) shows that the assumption of correct specification imposes potentially strong restrictions on interference by ruling out endogenous peer effects. We study the causal interpretation of $\tau(t,t^{\prime})$ under weak restrictions on interference, where the effect of $\bm{D}$ on $Y_{i}$ is not summarized by the econometrician’s choice of exposure mapping.
The identification results in this section apply to the class of estimands satisfying the following condition. We discuss estimands outside of this class in § 4.
Assumption 2 (Exposure Mapping).
For any $i$ and $\bm{d}\in\{0,1\}^{n}$, $f(\bm{d}_{\mathcal{N}(i,K)})=t^{\prime}$ implies $\bm{d}_{\mathcal{N}(i,K)}=\bm{\delta}_{i}$ for some $\bm{\delta}_{i}\in\{0,1\}^{\lvert\mathcal{N}(i,K)\rvert}$.
This is satisfied by choices of exposure mapping $T_{i}$ and exposure mapping realizations $t^{\prime}$ such that the event $T_{i}=t^{\prime}$ fully pins down the treatment subvector corresponding to $i$’s $K$-neighborhood. We have in mind the following examples, which are among the most common choices of exposure mappings found in the literature.
Example 1 (DIM).
Let $f(\bm{d}_{\mathcal{N}(i,K)})=d_{i}$, $t=1$, $t^{\prime}=0$, and $\mathcal{M}_{n}=\{1,\ldots,n\}$. Then $\tau(t,t^{\prime})$ is the population difference in means studied in § 2. Notice that $f(\bm{d}_{\mathcal{N}(i,K)})$ is a $K$-neighborhood exposure mapping for $K=0$, and $f(\bm{d}_{\mathcal{N}(i,K)})=t^{\prime}$ is equivalent to $\bm{d}_{\mathcal{N}(i,K)}=d_{i}=0$. Hence, Assumption 2 is satisfied.
Example 2 (Any Treated Neighbor).
Let $A_{ij}$ be an indicator for whether units $i$ and $j$ are linked in $\bm{A}$. Consider the exposure mapping $f(\bm{d}_{\mathcal{N}(i,K)})=(d_{i},\bm{1}\{\sum_{j=1}^{n}A_{ij}d_{j}>0\})$, $d\in\{0,1\}$, $t=(d,1)$, and $t^{\prime}=(d,0)$. Then $\tau(t,t^{\prime})$ is the average “effect” of having at least one treated neighbor for the subpopulation of units with treatment $d$. To ensure that $f(\bm{D}_{\mathcal{N}(i,K)})=t$ occurs with positive probability, we can choose $\mathcal{M}_{n}$ to be the subset of units with at least one neighbor. Assumption 2 is satisfied because $f(\bm{d}_{\mathcal{N}(i,K)})$ is a 1-neighborhood exposure mapping, and $f(\bm{d}_{\mathcal{N}(i,K)})=t^{\prime}$ implies that $d_{i}=d$ and all neighbors are untreated, which pins down $\bm{d}_{\mathcal{N}(i,K)}$.
Example 3 (Treated Neighbor Count).
Let $\mathcal{M}_{n}$ be the subset of units with $\gamma$ neighbors, $f(\bm{d}_{\mathcal{N}(i,K)})=(d_{i},\sum_{j=1}^{n}A_{ij}d_{j})$, $d\in\{0,1\}$, $\gamma\geq\tau$, $t=(d,\tau)$, and $t^{\prime}=(d,0)$. Then $\tau(t,t^{\prime})$ is the average “effect” of having zero versus $\tau$ treated neighbors for the subpopulation of units with treatment $d$ and $\gamma$ neighbors. The subpopulation must be chosen such that $\gamma\geq\tau$ or else the event $f(\bm{D}_{\mathcal{N}(i,K)})=t$ would occur with zero probability. Notice that $f(\bm{d}_{\mathcal{N}(i,K)})$ is a 1-neighborhood exposure mapping, and $f(\bm{d}_{\mathcal{N}(i,K)})=t^{\prime}$ implies that $d_{i}=d$ and all neighbors are untreated, so Assumption 2 is satisfied.
Suppose we instead take $t^{\prime}=(d,\tau^{\prime})$ for $0<\tau^{\prime}<\tau$. Then $\tau(t,t^{\prime})$ is the average “effect” of going from $\tau^{\prime}>0$ to $\tau$ treated neighbors for the same subpopulation of units as above. This does not generally satisfy Assumption 2 because knowing $\sum_{j\in\mathcal{N}(i,K)}d_{j}=\tau^{\prime}$ may not pin down $\bm{d}_{\mathcal{N}(i,K)}$. We study this estimand in § 4.
These are examples of 0- or 1-neighborhood exposure mappings. We will see instances of $K$-neighborhood exposure mappings for $K>1$ in Example 5 below.
3.2 Sign Reversal and Preservation
In § 2.1, we show that $\tau(1,0)$ is generally not informative for unit-level treatment effects when $T_{i}=D_{i}$. We provide a similar example here for indirect or spillover effects. Consider the exposure mapping in the first paragraph of Example 3, and suppose $n=3$, $\gamma=1$, $d=1$, $\tau=1$, $\mathcal{C}_{i}=\varnothing$, and the network is given by the following figure.
123
Write $Y_{i}(d_{1},d_{2},d_{3})\equiv Y_{i}(\bm{d})=0$ for $\bm{d}=(d_{i})_{i=1}^{3}\in\{0,1\}^{3}$. Abbreviate $\sigma_{i}(x,y,z\mid a,b)\equiv{\bf P}(\bm{D}=(x,y,z)\mid T_{i}=(a,b))$, and set
$Y_{1}(1,0,0)=Y_{3}(0,0,1)=0$ $\sigma_{1}(1,0,0\mid 1,0)=\sigma_{3}(0,0,1\mid 1,0)=p_{1}$
$Y_{1}(1,1,0)=Y_{3}(0,1,1)=1$ $\sigma_{1}(1,1,0\mid 1,1)=\sigma_{3}(0,1,1\mid 1,1)=p_{2}$
$Y_{1}(1,0,1)=Y_{3}(1,0,1)=2$ $\sigma_{1}(1,0,1\mid 1,0)=\sigma_{3}(1,0,1\mid 1,0)=p_{3}$
$Y_{1}(1,1,1)=Y_{3}(1,1,1)=3$ $\sigma_{1}(1,1,1\mid 1,1)=\sigma_{3}(1,1,1\mid 1,1)=p_{4}$
and $Y_{i}(\bm{d})=0$ for all other cases. Once again this example can be scaled up to larger $n$ by considering a large sample of identical such triads. Since $\gamma=1$, the subpopulation is $\mathcal{M}_{n}=\{1,3\}$, and for all such units, the unit-level spillover effects on the treated (e.g. $Y_{1}(1,1,d_{3})-Y_{1}(1,0,d_{3})$) are homogeneous, equaling one. Consider the assignment mechanism that draws uniformly at random from the subset of assignment vectors with exactly two units treated. Then $(p_{1},p_{2},p_{3},p_{4})=(0,1,1,0)$, so $\tau(t,t^{\prime})=-1$.
General sign preservation (Definition 2) does not restrict the sign of $\tau(t,t^{\prime})$ in this example because neither $\min_{i}\min\mathcal{T}_{i}(1,0)\geq 0$ nor $\max_{i}\max\mathcal{T}_{i}(1,0)\leq 0$. In particular, both $Y_{1}(1,1,1)-Y_{1}(1,0,1)>0$ and $Y_{1}(1,1,0)-Y_{1}(1,0,1)<0$. A criterion better tailored to this choice of exposure mapping and ceteris peribus comparisons would require that, if all unit-level spillover effects on the treated are positive (negative), then the sign of the aggregate “effect” $\tau(t,t^{\prime})$ is positive (negative). [^3] This example concerns a spillover effect, but it is also straightforward to engineer a sign reversal for $\tau(t,t^{\prime})$ when $t=(1,1)$ and $t^{\prime}=(0,0)$, which would correspond to a total effect that combines both direct and spillover effects. This notion of total effect is distinct from that of § 2.3, which involves a mediated effect.
We next define a generic sign preservation criterion for $K$-neighborhood exposure mappings that rules out the reversal in the previous example.
Definition 3.
For $t,t^{\prime}\in\mathbb{R}^{d_{t}}$ and an exposure mapping $f$, define the comparison set
$\mathcal{T}_{i}^{*}(t,t^{\prime})=\big{\{}Y_{i}(\bm{d})-Y_{i}(\bm{d}^{\prime})\colon\bm{d},\bm{d}^{\prime}\in\{0,1\}^{n},\bm{d}_{-\mathcal{N}(i,K)}=\bm{d}_{-\mathcal{N}(i,K)}^{\prime},\\ f(\bm{d}_{\mathcal{N}(i,K)})=t,f(\bm{d}_{\mathcal{N}(i,K)}^{\prime})=t^{\prime}\big{\}}.$
The estimand $\tau(t,t^{\prime})$ defined by exposure mapping $f$ satisfies $K$-neighborhood sign preservation if $\min_{i}\min\mathcal{T}_{i}^{*}(t,t^{\prime})\geq 0$ implies $\tau(t,t^{\prime})\geq 0$ and $\max_{i}\max\mathcal{T}_{i}^{*}(t,t^{\prime})\leq 0$ implies $\tau(t,t^{\prime})\leq 0$.
The difference relative to general sign preservation is that $\mathcal{T}_{i}^{*}(t,t^{\prime})$ holds fixed $\bm{d}_{-\mathcal{N}(i,K)}$ in the unit-level contrasts, which is the treatment subvector not directly manipulated by variation in the exposure mapping. This makes the comparison similar to a partial effect, but the fact that the exposure mapping may manipulate the entire $K$-neighborhood treatment subvector is reminiscent of a total effect.
In the context of the previous sign reversal example, $\mathcal{T}_{1}^{*}(t,t^{\prime})=\{Y_{1}(1,1,d_{3})-Y_{1}(1,0,d_{3})\colon d_{3}\in\{0,1\}\}$ and similarly for unit 3. These sets only contain unit-level spillover effects on the treated, which are all positive in the example, so Definition 3 rules out the sign reversal. This is to say that, if these spillover effects are the ceteris peribus comparisons of interest, then $K$-neighborhood sign preservation is the preferred criterion.
Let us specialize Definition 3 to each of the examples covered so far. First, if exposure mappings are correctly specified, then $K$-neighborhood sign preservation trivially holds since $Y_{i}(\bm{d})$ no longer varies with $\bm{d}_{-\mathcal{N}(i,K)}$. Second, for the estimand in Example 1, the criterion coincides with treatment sign preservation. Third, for the estimands in Example 2 and the first paragraph of Example 3, the comparison set simplifies to
$\mathcal{T}_{i}^{*}(t,t^{\prime})=\big{\{}Y_{i}(\bm{d}_{\mathcal{N}(i,1)},\bm{d}_{-\mathcal{N}(i,1)})-Y_{i}(\bm{\delta}_{i},\bm{d}_{-\mathcal{N}(i,1)})\colon\bm{d},\in\{0,1\}^{n},f(\bm{d}_{\mathcal{N}(i,1)})=t\big{\}},$
where $\bm{\delta}_{i}$ specifies that the ego is assigned treatment $d$ while all neighbors are untreated. The aggregate contrast $\tau(t,t^{\prime})$ compares this scenario with $T_{i}=t$, where the ego’s treatment remains fixed at $d$, but some neighbors become treated. The comparison set $\mathcal{T}_{i}^{*}(t,t^{\prime})$ is therefore the set of analogous unit-level comparisons, and Definition 3 demands a minimal level of consistency with the aggregate comparison.
The remainder of this section provides sufficient conditions for $K$-neighborhood sign preservation when the estimand satisfies Assumption 2. The second paragraph of Example 3 highlights an important estimand that fails to satisfy the assumption. In § 4, we will show that $K$-neighborhood sign preservation is actually too weak to rule out unpalatable sign reversals relevant to this estimand, which motivates the definition of a stronger criterion better tailored to it.
3.3 Unrestricted Potential Outcomes
We first show that $K$-neighborhood sign preservation can be satisfied under restrictions on selection into treatment while imposing no assumptions on potential outcomes. The main assumption is the following.
Assumption 3 (CI Selection).
For any $i$, $\bm{d}\in\{0,1\}^{n}$, and $s\in\{t,t^{\prime}\}$,
${\bf P}(\bm{D}=\bm{d}\mid T_{i}=s,\mathcal{C}_{i})\\ ={\bf P}(\bm{D}_{\mathcal{N}(i,K)}=\bm{d}_{\mathcal{N}(i,K)}\mid T_{i}=s,\mathcal{C}_{i})\,{\bf P}(\bm{D}_{-\mathcal{N}(i,K)}=\bm{d}_{-\mathcal{N}(i,K)}\mid\mathcal{C}_{i}).$
This imposes a form of conditional independence in treatment assignments. Whether the assumption holds depends on both on selection and the choice of controls $\mathcal{C}_{i}$. The simplest example is the setting of § 2, where $\bm{D}$ is randomly assigned independently across units. Here $K=0$, so $\bm{D}_{\mathcal{N}(i,K)}=D_{i}$, and Assumption 3 is satisfied with $\mathcal{C}_{i}=\varnothing$.
We first present the main result and then discuss primitive conditions for Assumption 3 in observational settings.
Theorem 1.
Under Assumptions 13, $\tau(t,t^{\prime})=\tau^{*}(t,t^{\prime})$ for
$\tau^{*}(t,t^{\prime})\equiv\frac{1}{m_{n}}\sum_{i\in\mathcal{M}_{n}}\bigg{(}\sum_{\bm{d}\in\{0,1\}^{n}}{\bf E}\big{[}Y_{i}(\bm{d})-Y_{i}(\bm{\delta}_{i},\bm{d}_{-\mathcal{N}(i,K)})\mid\mathcal{C}_{i}\big{]}\,\\ \times{\bf P}(\bm{D}=\bm{d}\mid T_{i}=t,\mathcal{C}_{i})\bigg{)},$ (3)
which satisfies $K$-neighborhood sign preservation.
Proof. See § A. Lemma A.1 uses Assumptions 1 and 2 to write $\tau(t,t^{\prime})$ as $\tau^{*}(t,t^{\prime})$ plus a bias term. The bias is zero under Assumption 3.
We study the economic content of Assumption 3 under the following model of interference proposed by Leung and Loupos (2024). For any $n$ and $i\in\{1,\ldots,n\}$, let
$Y_{i}=g_{n}(i,\bm{D},\bm{X},\bm{A},\bm{\varepsilon})\quad\text{and}\quad D_{i}=h_{n}(i,\bm{X},\bm{A},\bm{\nu}),$ (4)
where $\bm{X}=(X_{i})_{i=1}^{n}$ is an array of unit-level observables, $\bm{\varepsilon}=(\varepsilon_{i})_{i=1}^{n}$ is an array of outcome unobservables, $\bm{\nu}=(\nu_{i})_{i=1}^{n}$ is an array of selection unobservables, and $\{(g_{n},h_{n})\}_{n\in\mathbb{N}}$ a sequence of function pairs such that each $g_{n}(\cdot)$ has range $\mathbb{R}$ and $h_{n}(\cdot)$ has range $\{0,1\}$.
The timing of the model is (a) structural primitives $(\bm{A},\bm{X},\bm{\varepsilon},\bm{\nu})$ are realized; (b) units select into treatment according to a (possible) simultaneous-equations model with reduced form $h_{n}(\cdot)$; and (c) outcomes are realized according to a simultaneous-equations model with reduced form $g_{n}(\cdot)$. This allows for a complex form of interference in both the outcome and selection stages, for example interference induced by endogenous peer effects or strategic interactions more generally.
Example 4.
Suppose selection into treatment is determined by a game of complete information in which units take up treatment to maximize utility:
$D_{i}=\bm{1}\big{\{}U_{i}(\bm{D}_{-i},\bm{X},\bm{A},\bm{\nu})>0\big{\}},$
where the net utility $U_{i}(\ldots)$ from taking up treatment depends on peer take up $\bm{D}_{-i}$ (Bajari et al., 2010a; Balat and Han, 2023). The equilibrium selection mechanism is a reduced-form mapping from the primitives $(\bm{X},\bm{A},\bm{\nu})$ to an equilibrium outcome $\bm{D}$, which characterizes $D_{i}$ as a function $h_{n}(i,\bm{X},\bm{A},\bm{\nu})$.
In a game of incomplete information, $\nu_{i}$ is typically modeled as private information. Units decide whether to take up treatment to maximize expected utility, so that
$D_{i}=\bm{1}\big{\{}{\bf E}[U_{i}(\bm{D}_{-i},\bm{X},\bm{A},\bm{\nu})\mid\bm{X},\bm{A},\nu_{i}]>0\big{\}},$ (5)
where $(\bm{X},\bm{A},\nu_{i})$ is the information set of unit $i$ (Bajari et al., 2010b; Xu, 2018). This characterizes $D_{i}$ as a function $h_{n}(i,\bm{X},\bm{A},\nu_{i})$.
Under model (4), potential outcomes are given by
$Y_{i}(\bm{d})=g_{n}(i,\bm{d},\bm{X},\bm{A},\bm{\varepsilon}).$
By inspection of the selection equation, the entirety of $(\bm{X},\bm{A})$ is a potential source of confounding, so we take the controls to be
$\mathcal{C}_{i}=(\bm{X},\bm{A})\quad\text{for all }i.$ (6)
Then unconfoundedness as formulated in Assumption 1 is equivalent to $\bm{\varepsilon}\perp\!\!\!\perp\bm{\nu}\mid(\bm{X},\bm{A})$.
Proposition 1.
Suppose there exists a function $h_{n}(\cdot)$ such that $D_{i}=h_{n}(i,\bm{X},\bm{A},\nu_{i})$ for all $i$. Further suppose controls are chosen according to (6), and $\{\nu_{j}\}_{j=1}^{n}$ is independently distributed conditional on $\mathcal{C}_{i}$ for all $i$. Then $\{D_{j}\}_{j=1}^{n}$ is independently distributed conditional on $\mathcal{C}_{i}$ for all $i$, so Assumption 3 holds.
Proof. It is immediate from the assumptions that $\{D_{j}\}_{j=1}^{n}$ is independently distributed conditional on $\mathcal{C}_{i}$ for all $i$. This implies $\bm{D}_{-\mathcal{N}(i,K)}\perp\!\!\!\perp\bm{D}_{\mathcal{N}(i,K)}\mid\bm{X},\bm{A}$, so
${\bf P}(\bm{D}=\bm{d}\mid\underbrace{f(\bm{D}_{\mathcal{N}(i,K)})}_{T_{i}}=s,\underbrace{\bm{X},\bm{A}}_{\mathcal{C}_{i}})\\ ={\bf P}(\bm{D}_{\mathcal{N}(i,K)}=\bm{d}_{\mathcal{N}(i,K)}\mid f(\bm{D}_{\mathcal{N}(i,K)})=s,\bm{X},\bm{A})\\ \times{\bf P}(\bm{D}_{-\mathcal{N}(i,K)}=\bm{d}_{-\mathcal{N}(i,K)}\mid\bm{X},\bm{A}),$
which verifies Assumption 3.
The result provides primitive conditions for Assumption 3. Example 4 shows that the distinction between the models $D_{i}=h_{n}(i,\bm{X},\bm{A},\nu_{i})$ and $D_{i}=h_{n}(i,\bm{X},\bm{A},\bm{\nu})$ has economic content; the former is compatible with a game of incomplete information but generally not complete information, unlike the latter.
Combined with the previous theorems, the implication is that $K$-neighborhood sign preservation holds under a model with unrestricted interference in the outcome stage if (a) treatment selection follows a game of incomplete information, (b) private information $\nu_{i}$ is conditionally independent across units, and (c) we control for the entirety of $(\bm{X},\bm{A})$. Stronger versions of (b) are standard in structural analyses of model (5) (e.g. Lin and Vella, 2021; Lin and Xu, 2017; Xu, 2018). The upshot is that, for nonparametric selection models with incomplete information, we can establish $K$-neighborhood sign preservation without imposing any assumptions on the magnitude of strategic interactions in either the outcome or treatment selection stage.
These results purely concern identification, but estimation of $\tau(t,t^{\prime})$ is challenging due to high-dimensional network confounding (c). Leung and Loupos (2024) study nonparametric estimation and inference in this setting using graph neural networks.
3.4 Unrestricted Selection into Treatment
We next study the causal content of $\tau(t,t^{\prime})$ without imposing restrictions on selection into treatment (other than Assumption 1). If the exposure mapping is correctly specified, then $\tau(t,t^{\prime})$ immediately obtains a causal interpretation and satisfies sign preservation. Since correct specification can be a demanding requirement, we study to what extent it can be relaxed.
We assume outcomes are realized according to the model in (4). For any $S\subseteq\mathcal{N}_{n}$, let $\bm{A}_{S}=(A_{ij})_{i,j\in S}$ denote the submatrix of $\bm{A}$ restricted to $S$, $\bm{D}_{S}=(D_{i})_{i\in S}$, and similarly define $\bm{X}_{S}$ and other such submatrices. If a $K$-neighborhood exposure mapping is correctly specified, this implies that units further than distance $K$ do not interfere with the ego. We relax this to approximate neighborhood interference (ANI, Leung and Loupos, 2024, Assumption 2), which allows interference to decay more smoothly with path distance in the following sense.
Assumption 4 (ANI).
There exists a sequence of functions $\{\gamma_{n}(\cdot)\}_{n\in\mathbb{N}}$ with $\gamma_{n}\colon\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$ such that $\sup_{n\in\mathbb{N}}\gamma_{n}(s)\stackrel{{\scriptstyle s\rightarrow\infty}}{{\longrightarrow}}0$ and, for any $n\in\mathbb{N}$,
$\max_{i}{\bf E}\big{[}\lvert g_{n}(i,\bm{D},\bm{X},\bm{A},\bm{\varepsilon})\\ -g_{\lvert\mathcal{N}(i,s)\rvert}(i,\bm{D}_{\mathcal{N}(i,s)},\bm{X}_{\mathcal{N}(i,s)},\bm{A}_{\mathcal{N}(i,s)},\bm{\varepsilon}_{\mathcal{N}(i,s)})\rvert\mid\bm{D},\bm{X},\bm{A}\big{]}\leq\gamma_{n}(s)\quad\text{a.s.}$
We interpret $g_{\lvert\mathcal{N}(i,s)\rvert}(i,\dots)$ as $i$’s outcome under a counterfactual “$s$-neighborhood model” in which model primitives and treatments are fixed at their realizations, units external to the $s$-neighborhood are excluded from the model, and the remaining units interact according to the process $g_{\lvert\mathcal{N}(i,s)\rvert}(\cdot)$. The assumption bounds the difference between $i$’s realized outcome $g_{n}(i,\dots)$ and counterfactual $s$-neighborhood outcome $g_{\lvert\mathcal{N}(i,s)\rvert}(i,\dots)$ by $\gamma_{n}(s)$, which is required to decay with the radius $s$. This formalizes the idea that units increasingly distant from $i$ have increasingly less effect on $Y_{i}$.
[…]
Theorem 2.
Under Assumptions 1, 2, and 4, $\lvert\tau(t,t^{\prime})-\tau^{*}(t,t)^{\prime}\rvert\leq\gamma_{n}(K)$ a.s.
Proof. See § A. Lemma A.1 uses Assumptions 1 and 2 to write $\tau(t,t^{\prime})$ as $\tau^{*}(t,t^{\prime})$ defined in (3) plus a bias term. The bound on the bias is due to Assumption 4.
This result provides an “approximate” causal interpretation for $\tau(t,t^{\prime})$; exposure mappings capturing interference in a wider neighborhood around the ego result in estimands that are closer to a quantity satisfying $K$-neighborhood sign preservation. We next provide examples of exposure mappings that may be of interest when $K>1$.
Example 5 (Subnetwork Exposure).
Suppose $(\bm{\delta},\bm{a}),(\bm{\delta}^{\prime},\bm{a})$ are elements in the support of $(\bm{D}_{\mathcal{N}(i,K)},\bm{A}_{\mathcal{N}(i,K)})$, which we assume does not depend on $i$ [^4] This holds under weak exchangeability conditions, e.g. §3.3 of Leung and Loupos (2024). Define
$T_{i}=\left\{\begin{array}[]{ll}1&\text{if }(\bm{D}_{\mathcal{N}(i,K)},\bm{A}_{\mathcal{N}(i,K)})\cong(\bm{\delta},\bm{a})\\ 0&\text{if }(\bm{D}_{\mathcal{N}(i,K)},\bm{A}_{\mathcal{N}(i,K)})\cong(\bm{\delta}^{\prime},\bm{a})\end{array}\right.$
and $\mathcal{M}_{n}=\{i\colon\bm{A}_{\mathcal{N}(i,K)}\cong\bm{a}\}$, where $\cong$ denotes isomorphism. [^5] Formally, define a permutation as a bijection with domain $\{1,\ldots,n\}$. Abusing notation, write $\pi(\bm{D})=(D_{\pi(i)})_{i=1}^{n}$ and similarly $\pi(\bm{A})=(A_{\pi(i)\pi(j)})_{i,j}$, which permutes the rows and columns of the matrix $\bm{A}$. If there exists a permutation $\pi$ such that $(\bm{D}_{\mathcal{N}(i,K)},\bm{A}_{\mathcal{N}(i,K)})=(\pi(\bm{\delta}),\pi(\bm{a}))$, then we write $(\bm{D}_{\mathcal{N}(i,K)},\bm{A}_{\mathcal{N}(i,K)})\cong(\bm{\delta},\bm{a})$. Then $\tau(1,0)$ is the “effect” of changing a unit’s $K$-neighborhood treatment configuration from $\bm{\delta}^{\prime}$ to $\bm{\delta}$ for the subpopulation of units with $K$-neighborhood subnetwork $\bm{a}$. This is essentially the estimand studied in §4.3 of Auerbach and Tabord-Meehan (2023). Note that it satisfies Assumption 2.
The econometrician faces a bias-variance trade-off when selecting from a menu of $K$-neighborhood exposure mappings with different values of $K$. While bias shrinks with $K$ in the sense of Theorem 2, overlap quickly becomes limited for $K>1$. For instance, in Example 5, the proportion of units with a particular $K$-neighborhood treatment configuration can be very small in practice. For larger values of $K$, neighborhood subnetworks are extremely heterogeneous across units, and it can be difficult to find multiple units with identical such neighborhoods.
On the other hand, Examples 13 are only 0- and 1-neighborhood exposure mappings, so while they tend to have greater overlap in practice, the bias in Theorem 2 is relatively large, so the result reveals little about their causal interpretations. If we wish to justify the use of these estimands without imposing restrictions on selection into treatment as in § 3.3, we require stronger restrictions on interference, such as the following.
Proposition 2.
Suppose there exists $K^{\prime}\leq K$ such that potential outcomes only depend on the $K^{\prime}$-neighborhood treatment vector in that
$Y_{i}(\bm{d})=Y_{i}(\bm{d}^{\prime})\quad\text{for all}\quad\bm{d},\bm{d}^{\prime}\in\{0,1\}^{n}\colon\bm{d}_{\mathcal{N}(i,K^{\prime})}=\bm{d}_{\mathcal{N}(i,K^{\prime})}^{\prime}.$ (7)
Under Assumption 2, $\tau(t,t^{\prime})=\tau^{*}(t,t^{\prime})$, which satisfies $K$-neighborhood sign preservation.
[…]
Assumption (7) disallows interference from units beyond the ego’s $K^{\prime}$-neighborhood. This is a strong restriction, though it is weaker than correctly specified exposure mappings since it does not require $T_{i}$ to fully summarize $K$-neighborhood interference. For example, it could be that $f(\bm{d}_{\mathcal{N}(i,K)})=\max_{j\in\mathcal{N}(i,1)}d_{j}$ while $Y_{i}(\bm{d})$ depends on $\bm{d}$ through $(d_{i},\sum_{j=1}^{n}A_{ij}d_{j})$.
4 Ordered K-Neighborhood Sign Preservation
The results in § 3 apply to estimands satisfying Assumption 2, which include all but one of the estimands in Examples 13 and 5. In this section, we study the exception, which is the treated neighbor count estimand in the second paragraph of Example 3. This compares units with $\tau$ versus $\tau^{\prime}>0$ treated neighbors when the ego is assigned treatment $d$ and has $\gamma$ neighbors for $\gamma\geq\tau>\tau^{\prime}$. The next example shows that $K$-neighborhood sign preservation is too weak to rule out unpalatable sign reversals for this estimand.
4.1 Sign Reversal and Preservation
Consider a fully connected network with $n=4$, and (in the notation of Example 3) let $\gamma=3$, $d=1$, $\tau=2$, and $\tau^{\prime}=1$, so $\tau(t,t^{\prime})$ is the average “effect” of having two versus one neighbors treated for the subpopulation of treated units. Set
$Y_{1}(1,1,1,0)=1$ $Y_{1}(1,1,0,0)=0$
$Y_{1}(1,1,0,1)=2$ $Y_{1}(1,0,1,0)=1$
$Y_{1}(1,0,1,1)=3$ $Y_{1}(1,0,0,1)=2$
and $Y_{i}(\bm{d})=0$ for all other assignment vectors and units. Then $\tau(t,t^{\prime})={\bf E}[Y_{1}\mid T_{1}=t]-{\bf E}[Y_{1}\mid T_{1}=t^{\prime}]$. Observe that the unit-level spillover effect of going from one to two treated neighbors is weakly positive for all units, for example $Y_{1}(1,1,1,0)-Y(1,1,0,0)=1$ and $Y_{1}(1,1,0,1)-Y(1,1,0,0)=2$. Yet the aggregate comparison can be made negative by choosing ${\bf P}(\bm{D}=(1,1,1,0)\mid T_{1}=(1,2),\mathcal{C}_{1})={\bf P}(\bm{D}=(1,0,0,1)\mid T_{1}=(1,1),\mathcal{C}_{1})=1$, which results in $\tau(t,t^{\prime})=-1$.
As with the previous examples, the sign reversal occurs because we have a misspecified exposure mapping and dependent assignment mechanism. Since the network is fully connected, $K$-neighborhood sign preservation is equivalent to general sign preservation, and neither $\min_{i}\min\mathcal{T}_{i}^{*}(t,t^{\prime})\geq 0$ nor $\max_{i}\max\mathcal{T}_{i}^{*}(t,t^{\prime})\leq 0$ since $Y_{1}(1,1,1,0)-Y_{1}(1,1,0,0)>0$ and $Y_{1}(1,1,1,0)-Y_{1}(1,0,0,1)<0$. Hence neither criteria imposes restrictions on the sign of $\tau(t,t^{\prime})$.
We next state a stronger criterion that rules out the reversal.
Definition 4.
For $t,t^{\prime}\in\mathbb{R}^{d_{t}}$ and an exposure mapping $f$, define the comparison set
$\mathcal{T}_{i}^{o}(t,t^{\prime})=\big{\{}Y_{i}(\bm{d})-Y_{i}(\bm{d}^{\prime})\colon\bm{d},\bm{d}^{\prime}\in\{0,1\}^{n},\bm{d}_{-\mathcal{N}(i,K)}=\bm{d}_{-\mathcal{N}(i,K)}^{\prime}\\ \bm{d}_{\mathcal{N}(i,K)}\geq\bm{d}_{\mathcal{N}(i,K)}^{\prime},f(\bm{d}_{\mathcal{N}(i,K)})=t,f(\bm{d}_{\mathcal{N}(i,K)}^{\prime})=t^{\prime}\big{\}}.$
The estimand $\tau(t,t^{\prime})$ defined by exposure mapping $f$ satisfies ordered $K$-neighborhood sign preservation if $\min_{i}\min\mathcal{T}_{i}^{o}(t,t^{\prime})\geq 0$ implies $\tau(t,t^{\prime})\geq 0$ and $\max_{i}\max\mathcal{T}_{i}^{o}(t,t^{\prime})\leq 0$ implies $\tau(t,t^{\prime})\leq 0$.
The difference relative to $K$-neighborhood sign preservation is the added constraint $\bm{d}_{\mathcal{N}(i,K)}\geq\bm{d}_{\mathcal{N}(i,K)}^{\prime}$ (in the usual partial order). This ensures that the ceteris peribus comparisons in $\mathcal{T}_{i}^{o}(t,t^{\prime})$ only involve increases in the number of treated $K$-neighbors, holding all else equal. This criterion directly rules out the sign reversal in § 4.1 since the comparison set only contains contrasts involving partially ordered assignment vectors, one with two treated neighbors and the other with one. Hence, if these unit-level spillover effects are the ceteris peribus comparisons of interest, then Definition 4 is the preferred criterion.
Whether ordered $K$-neighborhood sign preservation is preferable to its unordered version, or whether some other criterion is more appropriate, depends on the exposure mapping and ceteris peribus comparisons of interest.
  • For the difference in means (Example 1) and any treated neighbor (Example 2) estimands, the ordered and unordered criteria coincide, and Theorems 1 and 2 provide sufficient conditions under which they are satisfied.
  • For the treated neighbor count estimand (Example 3) with $\tau^{\prime}>0$, the sign reversal example above provides a rationale for preferring ordered $K$-neighborhood sign preservation, and the next subsection discusses sufficient conditions. If $\tau^{\prime}=0$, then the two criteria coincide, and Theorems 1 and 2 provide sufficient conditions.
  • For the subnetwork exposure estimand (Example 5), if $\bm{\delta}$ and $\bm{\delta}^{\prime}$ are partially ordered, then the two criteria coincide. If they are not, then the comparison set $\mathcal{T}_{i}^{o}(t,t^{\prime})$ is empty, so the ordered criterion is vacuous, unlike the unordered version.
4.2 Sufficient Conditions
In § 3, we discuss two routes for establishing $K$-neighborhood sign preservation. In § 3.3, we leave potential outcomes unrestricted but impose assumptions on selection into treatment. In § 3.4, we take the opposite approach. We first observe that, when it comes to ordered $K$-neighborhood sign preservation, the second approach is not fruitful. The reason is that the sign reversal example in § 4.1 satisfies the restriction (7) on interference, demonstrating that ordered $K$-neighborhood sign preservation fails to hold under a slightly weaker assumption than correct specification. Put another way, this illustrates the necessity of Assumption 2 for the second route. The next theorem takes the first route.
Theorem 3.
[…]
Proof. See § A.
The assumption of i.i.d. treatments is obviously stronger than Assumption 3 and only plausible in experimental settings. We conjecture that this can be relaxed to the requirement that $\{D_{j}\}_{j=1}^{n}$ is independently distributed conditional on $\mathcal{C}_{i}$ for all $i$ (see Proposition 1 for sufficient conditions).
Under SUTVA, the typical argument for showing that an estimand has a causal interpretation is to rewrite it as a non-negatively weighted average of unit-level partial effects (see the references in § 1). This is the method of proof for Theorems 1 and 2 above. Assumption 2 is critical for this purpose since it enables us to compare the potential outcome under every treatment vector in the support of ${\bf P}(\bm{D}=\bm{d}\mid T_{i}=t,\mathcal{C}_{i})$ to a single potential outcome $Y_{i}(\bm{\delta}_{i},\bm{d}_{-\mathcal{N}(i,K)})$, resulting in (3).
Without Assumption 2, it is unclear whether $\tau(t,t^{\prime})$ has an analogous representation since there are generally multiple assignment vectors $\bm{D}$ for which $T_{i}=t$ and multiple for which $T_{i}=t^{\prime}$. The proof of Theorem 3 therefore relies on a different strategy. We construct a coupling $Y_{i}(\bm{D}_{\tau})\stackrel{{\scriptstyle a.s.}}{{\geq}}Y_{i}(\bm{D}_{\tau^{\prime}})$ for each $i$ such that $\bm{D}_{\tau}$ has distribution ${\bf P}(\bm{D}=\bm{d}\mid T_{i}=\tau,\mathcal{C}_{i})$ and $\bm{D}_{\tau^{\prime}}$ has distribution ${\bf P}(\bm{D}=\bm{d}\mid T_{i}=\tau^{\prime},\mathcal{C}_{i})$. Under this coupling, $\min\mathcal{T}_{i}^{o}(t,t^{\prime})\geq 0$ implies ${\bf E}[Y_{i}\mid T_{i}=t,\mathcal{C}_{i}]-{\bf E}[Y_{i}\mid T_{i}=t^{\prime},\mathcal{C}_{i}]\geq 0$. Hence, $\min_{i}\min\mathcal{T}_{i}^{o}(t,t^{\prime})\geq 0$ implies $\tau(t,t^{\prime})$.
5 Conclusion
Researchers commonly utilize exposure mappings to define estimands for interference. We demonstrate that the aggregate statistical comparison made by these estimands, such as contrasting the average outcomes of treated and untreated units, can have the opposite sign of all analogous unit-level ceteris peribus comparisons, even under standard identification conditions. Therefore, without further assumptions, the estimands are not informative for the unit-level comparisons.
We discuss a variety of sign preservation criteria for causal interpretability and illustrate how the choice of criterion depends on the exposure mapping and ceteris peribus comparisons of interest. To satisfy preferred criteria, it is necessary to impose restrictions on interference either for potential outcomes or selection into treatment. This mirrors the Wald estimand for instrumental variables, which obtains a causal interpretation under restrictions on heterogeneity either for potential outcomes or selection into treatment. We exhibit a nonparametric model under which it is possible to satisfy a preferred sign preservation criterion without restricting potential outcomes or the magnitude of peer effects in treatment adoption.
Appendix A Proofs
Lemma A.1.
Under Assumptions 1 and 2, $\tau(t,t^{\prime})=\tau^{*}(t,t^{\prime})+\mathcal{R}_{n}$, where $\tau^{*}(t,t^{\prime})$ is defined in (3) and
$\mathcal{R}_{n}=\frac{1}{m_{n}}\sum_{i\in\mathcal{M}_{n}}\bigg{(}\sum_{\bm{d}\in\{0,1\}^{n}}{\bf E}\big{[}Y_{i}(\bm{\delta}_{i},\bm{d}_{-\mathcal{N}(i,K)})\mid\mathcal{C}_{i}\big{]}\\ \times\,\big{(}{\bf P}(\bm{D}=\bm{d}\mid T_{i}=t,\mathcal{C}_{i})-{\bf P}(\bm{D}=\bm{d}\mid T_{i}=t^{\prime},\mathcal{C}_{i})\big{)}\bigg{)}.$
Proof. By Assumption 1,
$\tau(t,t^{\prime})=\frac{1}{m_{n}}\sum_{i\in\mathcal{M}_{n}}\bigg{(}\sum_{\bm{d}\in\{0,1\}^{n}}{\bf E}[Y_{i}(\bm{d})\mid\mathcal{C}_{i}]{\bf P}(\bm{D}=\bm{d}\mid T_{i}=t,\mathcal{C}_{i})\\ -\sum_{\bm{d}^{\prime}\in\{0,1\}^{n}}{\bf E}[Y_{i}(\bm{d}^{\prime})\mid\mathcal{C}_{i}]{\bf P}(\bm{D}=\bm{d}^{\prime}\mid T_{i}=t^{\prime},\mathcal{C}_{i})\bigg{)}.$ (8)
Under Assumption 2,
(8)$=\frac{1}{m_{n}}\sum_{i\in\mathcal{M}_{n}}\bigg{(}\sum_{\bm{d}\in\{0,1\}^{n}}{\bf E}\big{[}Y_{i}(\bm{d})\mid\mathcal{C}_{i}\big{]}\,{\bf P}(\bm{D}=\bm{d}\mid T_{i}=t,\mathcal{C}_{i})\\ -\sum_{\bm{d}^{\prime}\in\{0,1\}^{n}}{\bf E}\big{[}Y_{i}(\bm{\delta}_{i},\bm{d}_{-\mathcal{N}(i,K)}^{\prime})\mid\mathcal{C}_{i}\big{]}\,{\bf P}(\bm{D}=\bm{d}^{\prime}\mid T_{i}=t^{\prime},\mathcal{C}_{i})\bigg{)}.$
Add and subtract
$\frac{1}{m_{n}}\sum_{i\in\mathcal{M}_{n}}\sum_{\bm{d}\in\{0,1\}^{n}}{\bf E}\big{[}Y_{i}(\bm{\delta}_{i},\bm{d}_{-\mathcal{N}(i,K)})\mid\mathcal{C}_{i}\big{]}\,{\bf P}(\bm{D}=\bm{d}\mid T_{i}=t,\mathcal{C}_{i})$
from the right-hand side to obtain $\tau^{*}(t,t^{\prime})+\mathcal{R}_{n}$.
Proof of Theorem 1. By Lemma A.1, $\tau(t,t^{\prime})=\tau^{*}(t,t^{\prime})+\mathcal{R}_{n}$. By Assumption 3,
$\mathcal{R}_{n}=\frac{1}{m_{n}}\sum_{i\in\mathcal{M}_{n}}\bigg{(}\sum_{\bm{d}_{-\mathcal{N}(i,K)}}{\bf E}\big{[}Y_{i}((\bm{\delta}_{i},\bm{d}_{-\mathcal{N}(i,K)})\mid\mathcal{C}_{i}\big{]}{\bf P}(\bm{D}_{-\mathcal{N}(i,K)}=\bm{d}_{-\mathcal{N}(i,K)}\mid\mathcal{C}_{i})\\ \times\,\bigg{(}\sum_{\bm{d}_{\mathcal{N}(i,K)}}{\bf P}(\bm{D}_{\mathcal{N}(i,K)}=\bm{d}_{\mathcal{N}(i,K)}\mid T_{i}=t,\mathcal{C}_{i})\\ -\sum_{\bm{d}_{\mathcal{N}(i,K)}}{\bf P}(\bm{D}_{\mathcal{N}(i,K)}=\bm{d}_{\mathcal{N}(i,K)}\mid T_{i}=t^{\prime},\mathcal{C}_{i})\bigg{)}\bigg{)}.$
The sums in the last two lines equal one, so $\mathcal{R}_{n}=0$.
Proof of Theorem 2. By Lemma A.1, $\tau(t,t^{\prime})=\tau^{*}(t,t^{\prime})+\mathcal{R}_{n}$. Define
$\mathcal{R}_{n}^{*}=\frac{1}{m_{n}}\sum_{i\in\mathcal{M}_{n}}\bigg{(}\sum_{\bm{d}\in\{0,1\}^{n}}\big{(}{\bf E}\big{[}Y_{i}(\bm{\delta}_{i},\bm{d}_{-\mathcal{N}(i,K)})\mid\mathcal{C}_{i}\big{]}\\ -{\bf E}\big{[}g_{\lvert\mathcal{N}(i,K)\rvert}(i,\bm{\delta}_{i},\bm{X}_{\mathcal{N}(i,K)},\bm{A}_{\mathcal{N}(i,K)},\bm{\varepsilon}_{\mathcal{N}(i,K)})\mid\mathcal{C}_{i}\big{]}\big{)}\\ \times\,\big{(}{\bf P}(\bm{D}=\bm{d}\mid T_{i}=t,\mathcal{C}_{i})-{\bf P}(\bm{D}=\bm{d}\mid T_{i}=t^{\prime},\mathcal{C}_{i})\big{)}\bigg{)}.$
Then
$*.\mathcal{R}_{n}=\frac{1}{m_{n}}\sum_{i\in\mathcal{M}_{n}}{\bf E}\big{[}g_{\lvert\mathcal{N}(i,K)\rvert}(i,\bm{\delta}_{i},\bm{X}_{\mathcal{N}(i,K)},\bm{A}_{\mathcal{N}(i,K)},\bm{\varepsilon}_{\mathcal{N}(i,K)})\mid\mathcal{C}_{i}\big{]}\\ \times\underbrace{\sum_{\bm{d}\in\{0,1\}^{n}}\big{(}{\bf P}(\bm{D}=\bm{d}\mid T_{i}=t,\mathcal{C}_{i})-{\bf P}(\bm{D}=\bm{d}\mid T_{i}=t^{\prime},\mathcal{C}_{i})\big{)}}_{0}\,\,{+}\,\,\mathcal{R}_{n}^{*}.$
By Assumption 1, under the event $f(\bm{D}_{\mathcal{N}(i,K)})=t^{\prime}$,
${\bf E}\big{[}Y_{i}(\bm{\delta}_{i},\bm{d}_{-\mathcal{N}(i,K)})\mid\mathcal{C}_{i}\big{]}-{\bf E}\big{[}g_{\lvert\mathcal{N}(i,K)\rvert}(i,\bm{\delta}_{i},\bm{X}_{\mathcal{N}(i,K)},\bm{A}_{\mathcal{N}(i,K)},\bm{\varepsilon}_{\mathcal{N}(i,K)})\mid\mathcal{C}_{i}\big{]}\\ ={\bf E}[g_{n}(i,\bm{D},\bm{A},\bm{X},\bm{\varepsilon})-g_{\lvert\mathcal{N}(i,K)\rvert}(i,\bm{\delta}_{i},\bm{X}_{\mathcal{N}(i,K)},\bm{A}_{\mathcal{N}(i,K)},\bm{\varepsilon}_{\mathcal{N}(i,K)})\mid\bm{D},\bm{X},\bm{A}\big{]}.$
By Assumption 4, this is a.s. bounded above in absolute value by $\gamma_{n}(K)$. Hence, $\lvert\mathcal{R}_{n}\rvert\leq\gamma_{n}(K)$.
Proof of Theorem 3. Fix a unit $i$, and for any $s\in\{\tau,\tau^{\prime}\}$, let $\bm{D}_{s}$ denote a draw from the conditional distribution ${\bf P}(\bm{D}=\bm{d}\mid T_{i}=(d,s),\mathcal{C}_{i})$, a distribution constructed from the “true assignment mechanism” ${\bf P}(\bm{D}=\bm{d}\mid\mathcal{C}_{i})$. We will construct a coupling of $\bm{D}_{\tau}$ and $\bm{D}_{\tau^{\prime}}$ such that $\bm{D}_{\tau}\geq\bm{D}_{\tau^{\prime}}$ a.s., so that $\min\mathcal{T}_{i}^{o}(t,t^{\prime})\geq 0$ implies
$Y_{i}(\bm{D}_{\tau})\stackrel{{\scriptstyle a.s.}}{{\geq}}Y_{i}(\bm{D}_{\tau^{\prime}}),\quad\text{which implies}\\ {\bf E}[Y_{i}(\bm{D}_{\tau})-Y_{i}(\bm{D}_{\tau^{\prime}})]={\bf E}[Y_{i}\mid T_{i}=t,\mathcal{C}_{i}]-{\bf E}[Y_{i}\mid T_{i}=t^{\prime},\mathcal{C}_{i}]\geq 0.$
Hence if $\min_{i}\min\mathcal{T}_{i}^{o}(t,t^{\prime})\geq 0$, then $\tau(t,t^{\prime})\geq 0$. A similar argument can be used to show that $\min_{i}\min\mathcal{T}_{i}^{o}(t,t^{\prime})\leq 0$ implies $\tau(t,t^{\prime})\leq 0$, which completes the proof.
Abbreviate $p_{i}={\bf P}(D_{i}=1\mid\mathcal{C}_{i})$. Without loss generality, set $i=1$, and let units $12,\ldots\gamma+$ 1 be its neighbors. To construct $\bm{D}_{\tau^{\prime}}$, set unit 1’s treatment assignment to $d$, and for each $j\in\{\gamma+2,\ldots,n\}$ (the set of units excluding $i$’s 1-neighborhood), independently draw its treatment assignment from a Bernoulli$(p_{j})$ distribution. Consider an urn containing labels $\{2,\ldots,\gamma+1\}$. Randomly sample $\tau^{\prime}$ indices without replacement from the urn, and assign all units corresponding to the sampled indices to treatment. Call the treatment assignment vector resulting from this whole process $\bm{D}_{\tau^{\prime}}$. Next, randomly sample $\tau-\tau^{\prime}$ of the remaining indices in the urn without replacement. Construct $\bm{D}_{\tau}$ from $\bm{D}_{\tau^{\prime}}$ by switching the units corresponding to the newly sampled indices to treatment.
By construction, for each $s\in\{\tau,\tau^{\prime}\}$, $\bm{D}_{s}$ is an assignment vector on $\{0,1\}^{n}$ in which unit $i$ is assigned treatment $d$ and $s$ out of $\gamma$ neighbors are treated, and furthermore $\bm{D}_{\tau}\geq\bm{D}_{\tau^{\prime}}$. Hence if $\min\mathcal{T}_{i}^{o}(t,t^{\prime})\geq 0$, then $Y_{i}(\bm{D}_{\tau})\geq Y_{i}(\bm{D}_{\tau^{\prime}})$. It therefore remains to show that $\bm{D}_{s}$ is a draw from ${\bf P}(\bm{D}=\bm{d}\mid T_{i}=(d,s),\mathcal{C}_{i})$. We will only show this for $s=\tau$ as the other case is the same with $\tau$ replaced by $\tau^{\prime}$.
Let $N_{i}=\sum_{j=1}^{n}A_{ij}D_{j}$. Under the true assignment mechanism, treatments are independently distributed conditional on $\mathcal{C}_{i}$ for any $i$, and $T_{i}=t$ implies $D_{i}=d$. Then for any $\bm{d}\in\{0,1\}^{n}$ such that the $i$th component equals $d$ and $\sum_{i=2}^{\gamma+1}d_{i}=\tau$,
${\bf P}\big{(}\bm{D}=\bm{d}\mid T_{i}=t,\mathcal{C}_{i}\big{)}\\ ={\bf P}\big{(}(D_{i})_{i=2}^{\gamma+1}=(d_{i})_{i=2}^{\gamma+1}\mid N_{i}=\tau,\mathcal{C}_{i})\,{\bf P}\big{(}(D_{i})_{i=\gamma+2}^{n}=(d_{i})_{i=\gamma+2}^{n}\mid\mathcal{C}_{i}\big{)},$
and ${\bf P}(\bm{D}=\bm{d}\mid T_{i}=t,\mathcal{C}_{i})=0$ for any other $\bm{d}$. Since treatments are independently distributed conditional on $\mathcal{C}_{i}$,
${\bf P}\big{(}(D_{i})_{i=\gamma+2}^{n}=(d_{i})_{i=\gamma+2}^{n}\mid\mathcal{C}_{i}\big{)}=\frac{\prod_{i=\gamma+2}^{n}p_{i}^{d_{i}}(1-p_{i})^{1-d_{i}}}{\sum_{\{d_{j}\}_{j=\gamma+2}^{n}\subseteq\{0,1\}}\prod_{k=\gamma+2}^{n}p_{k}^{d_{k}}(1-p_{k})^{1-d_{k}}}.$
This expression is the same distribution as the subvectors of $\bm{D}_{\tau}$ and $\bm{D}_{\tau}$ restricted to $\{\gamma+2,\ldots,n\}$. Since $p_{i}=p$ for all $i$,
${\bf P}\big{(}(D_{i})_{i=2}^{\gamma+1}=(d_{i})_{i=2}^{\gamma+1}\mid N_{i}=\tau,\mathcal{C}_{i})\\ =\frac{\prod_{i=2}^{\gamma+1}p^{d_{i}}(1-p)^{1-d_{i}}}{\sum_{\begin{subarray}{c}\{d_{j}\}_{j=2}^{\gamma+1}\subseteq\{0,1\}\\ \sum_{j=2}^{\gamma+1}d_{j}=\tau\end{subarray}}\prod_{k=2}^{\gamma+1}p^{d_{k}}(1-p)^{1-d_{k}}}=\binom{\gamma}{\tau}^{-1}.$
The right-hand side equals the chance of randomly sampling $\tau$ out of $\gamma$ indices without replacement and therefore corresponds to the distribution of the subvector of $\bm{D}_{\tau}$ on $\{2,\ldots,\gamma+1\}$.